OpenStudy (anonymous):

Find an equation of the circle that satisfies the given conditions. Center (−10, 5); passes through (9, 3)

OpenStudy (anonymous):

use the standard form of the equation of a circle

OpenStudy (anonymous):

the standard form is: (x-a)^2+(y-b)^2=r^2; c^; (x+10)^2+(y-5)^2=sqrt(365) disregard the c^ in the middle there. my delete button is failing right now